# FiniteTimeBibliography.bib

### v1

@@ Appendix A Quantum Mazur-Suzuki inequality over finite times
Eq. (184) is nontrivial when the right hand side is larger than $A_{\mathrm{th}}^{2}$, which is the minimum value of the autocorrelator and corresponds to (global) thermalization. However, it only constrains thermalization over the same timescale (determined by $\Delta E$) in which the charge remains approximately conserved. For a noticeably broken conservation law (say, after a long timescale), $\delta Q_{k}^{2}[w_{2+}]$ is large, and consequently the second term within the absolute value that depends on $\delta Q_{k}^{2}[w_{2+}]$ weakens the bound, allowing the autocorrelator to decay further than with a perfect conservation law and potentially thermalize (as expected for systems with fewer conservation laws). For the question [DharMazur] of when this inequality can be saturated over finite timescales with a finite number of accessible charges $Q_{k}$, which are not necessarily energy projectors as in eigenstate thermalization [Eq. (12)], we note that this is possible when $[\hat{A}]_{\Delta E}$ has most of its overlap with the $\hat{Q}_{k}$ rather than the $\hat{J}_{k}$ by Eq. (177) and the completeness relation Eq. (176) with $\hat{O}=[\hat{A}]_{\Delta E}$. It remains to be seen if this criterion can be used more systematically to determine when the inequality may be saturated in different classes of systems. It would also be interesting to consider the implications of the exact inequality in Eq. (184) for quantum transport problems, e.g., in a finite but large number of qubits.
@@ Appendix B Thermalization in dual-unitary quantum circuits
Under these circumstances, it can be shown that all autocorrelators of traceless single-site observables vanish for all time steps $\lvert t\rvert<N$, provided each unitary $2$-qubit gate satisfies the property of “dual unitarity”, for the details of which we refer to Ref. [ProsenErgodic]. While the dynamics of specific (matrix product) initial states have been computed exactly for such circuits [PiroliDualUnitary], and $f(E_{1},E_{2})$ in the ETH ansatz for certain observables has been estimated from autocorrelators [FritzschProsen] assuming the form in Eq. (3), our intention here is to show how more general conclusions on thermalization may be drawn in a much simpler way. Quantitatively, all traceless single-qubit observables $\widehat{\delta a}_{i}$ (with $i=1$ to $N$ indexing the qubit) satisfy:
$\operatorname{Tr}[\widehat{\delta a}_{i}(t)\widehat{\delta a}_{i}(0)]=0,\text{for all }t:0<|t|<N.$ (185)
For intuition, we note that in the $N\to\infty$ thermodynamic limit, all autocorrelators thermalize at all (nonzero) finite times. Then, our results [informally: Eq. (11) and the discussion around (15); formally: Summary 5.4 and the Theorems therein with Sec. 7] imply the following: Given any basis of initial states $\lvert\psi_{k}\rangle$ (which may be the computational basis states, or more nontrivially any family of states obtained by acting with an arbitrary unitary circuit on the computational basis states), and every single-qubit observable $\hat{a}_{i}$, for every $\epsilon>0$ , there exists some $N_{0}\in\mathbb{N}$ and $T_{0}\in\mathbb{N}$ such that:
$\left\lvert\langle\psi_{k}(t)\rvert\ \hat{a}_{i}\ \lvert\psi_{k}(t)\rangle-\frac{1}{2}\operatorname{Tr}_{i}[\hat{a}_{i}]\right\rvert<\epsilon\text{ for almost all }t\in[-T,T],\text{ for any }T>T_{0},$ (186)
@@
While we have expressed Eq. (186) for single-qubit observables, larger local observables on, say, $N_{S}$ consecutive qubits must also have vanishing autocorrelators except at extremely short $t=O(N_{S})$ and extremely long $t=\Omega(N)$ times, by our understanding of Ref. [ProsenErgodic]. We therefore expect an analogue of Eq. (186) to hold for any traceless local observable spanning $N_{S}=o(N)$ consecutive sites.
Despite the strong thermalization behavior indicated by Eq. (186), it is known that some of these circuits show spectral correlations corresponding to dynamical non-ergodicity, e.g., Poisson spectral statistics [ClaeysErgodicCircuits]. These examples therefore also directly illustrate the logical separation between ergodic dynamics and the statistical mechanics of observables.
@@
$\hat{a}_{i}^{\text{co}}[t]=\hat{a}_{i+2t}.$ (187)
[…]
@@ Appendix C Energy-band thermalization $\implies$ eigenspaces usually thermalize
$\Sigma_{d}=\Sigma_{d_{s}}\oplus\Sigma_{\delta d}$ (188)
@@
$n(\Sigma_{\delta d})\equiv\left(\sum_{\mathcal{H}_{\delta}(E)\in\Sigma_{\delta d}}1\right)<\frac{\epsilon^{2}}{\lambda^{2}}d.$ (189)
@@
Let us consider the implications of Eq. (189). First, we note that for almost all eigenspaces to thermalize, the ratio $n(\Sigma_{\delta d})/n(\Sigma_{d})$ should become negligible:
$\frac{n(\Sigma_{\delta d})}{n(\Sigma_{d})}\ll 1.$ (190)
In analogy with “weak eigenstate thermalization” [BiroliWeakETH, MoriWeakETH, MoriETHreview], we will diagnose “weak eigenspace thermalization” if Eq. (190) is satisfied. By Proposition 3.3, this implies that thermalization o.a. occurs in arbitrary initial states supported on an energy shell $\Sigma_{d_{s}}$ that contains most of the eigenspaces in $\Sigma_{d}$. Determining weak eigenspace thermalization requires our accuracy for energy-band thermalization to be:
$\epsilon_{\text{weak}}\ll\lambda\ \sqrt{\frac{n(\Sigma_{d})}{d}}.$ (191)
@@
$\epsilon_{\text{weak}}>\frac{1}{N^{\alpha}},$ (192)
@@
$n(\Sigma_{d})>\frac{d}{N^{2\alpha}}.$ (193)
That we cannot show weak eigenspace thermalization with an accessible accuracy of energy-band thermalization without some constraint on the spectrum is why the title of this Appendix emphasizes the qualifier “usually”. Concretely, “usually” refers to systems satisfying Eq. (193); it implies that the spectrum must generally have a low degree of degeneracy. Nevertheless, we expect this condition to apply to a large class of systems, especially seeing that it allows $n(\Sigma_{d})\ll d$ (such as systems with a finite or at most weakly growing degree of degeneracy in each level, in the thermodynamic limit $d\to\infty$).
@@
$\epsilon<\frac{\lambda}{\sqrt{d}}$ (195)
in energy-band thermalization. This is consistent with Eq. (63). This constraint is independent of the energy spectrum, but as $\lambda<1$ for a nontrivial expression of eigenspace thermalization, Eq. (195) is not expected to be an accessible degree of accuracy (scaling as a power of $1/d$).
@@ D.1 Classical “eigenstate” thermalization without ergodicity
$=\sum_{k}\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}\pi_{\mathcal{P}_{k}}[\rho].$ (196)
@@
From Eq. (35), rewriting the left hand side in terms of the ergodic subsets $\mathcal{P}_{k}$ as in Eq. (196), we get
$\sum_{k}\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}\pi_{\mathcal{P}_{k}}[\rho_{A}]=\frac{\mu(A)}{\mu(\mathcal{P})}.$ (198)
@@
$\pi_{\mathcal{P}_{k}}[\rho_{A}]=\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(A)},$ (199)
using which Eq. (198) becomes (together with some manipulations on the right hand side using $\sum_{k}\mu(\mathcal{P}_{k})=\mu(\mathcal{P})$):
@@
$\implies\frac{1}{\mu(A)}\sum_{k}\mu(\mathcal{P}_{k})\frac{\mu(A\cap\mathcal{P}_{k})^{2}}{\mu(\mathcal{P}_{k})^{2}}$ $=\frac{1}{\mu(A)}\left[2\sum_{k}\mu(\mathcal{P}_{k})\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}\frac{\mu(A)}{\mu(\mathcal{P})}-\sum_{k}\mu(\mathcal{P}_{k})\frac{\mu(A)^{2}}{\mu(\mathcal{P})^{2}}\right].$ (200)
Collecting all the terms in Eq. (200), we get a weighted variance:
$\frac{1}{\mu(A)}\sum_{k}\mu(\mathcal{P}_{k})\left[\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}-\frac{\mu(A)}{\mu(\mathcal{P})}\right]^{2}=0.$ (201)
@@
$\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}-\frac{\mu(A)}{\mu(\mathcal{P})}=0.$ (202)
@@ D.2 Quantum eigenspace thermalization with degeneracies
D.2.1 Proposition 3.1: Quantum eigenstate thermalization implies thermalization o.a. given a nondegenerate spectrum [vonNeumannThermalization, srednicki1999eth]
@@
$\left\lvert\int\overline{\mathrm{d}t}\ \operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert\leq\left(\sum_{n:\lvert E_{n}\rangle\in\Sigma_{d}}\langle E_{n}\rvert\hat{\rho}\lvert E_{n}\rangle\right)\max_{n:\lvert E_{n}\rangle\in\Sigma_{d}}\left\lvert\langle E_{n}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert$ (203)
@@
Rewriting Eq. (50) to include the $-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}$ term, we get:
@@
$=\sum_{E\in\mathcal{E}}\operatorname{Tr}\left[\hat{\Pi}(E)\hat{\rho}\hat{\Pi}(E)\hat{\Pi}(E)\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}(E)\right],$ (204)
@@
$\hat{\Pi}(E)=\hat{\Pi}(E)^{2}=\sum_{n:\lvert E_{n}\rangle\in\mathcal{H}_{E}}\lvert E_{n}\rangle\langle E_{n}\rvert.$ (205)
Now applying the Cauchy-Schwarz inequality to each term in Eq. (204) with a given $E$, and the triangle inequality to consider each such term separately, we get
$\left\lvert\int\overline{\mathrm{d}t}\ \operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert$ $\leq\sum_{E\in\mathcal{E}}\sqrt{\operatorname{Tr}\left[\left\{\hat{\rho}\hat{\Pi}(E)\right\}^{2}\right]\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}(E)\right\}^{2}\right]}$ (206)
$\leq\epsilon\sum_{E\in\mathcal{E}}\sqrt{\operatorname{Tr}\left[\left\{\hat{\rho}\hat{\Pi}(E)\right\}^{2}\right]}.$ (207)
@@
$\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\rvert^{2}\leq\langle E_{n}\rvert\hat{\rho}\lvert E_{n}\rangle\langle E_{m}\rvert\hat{\rho}\lvert E_{m}\rangle,$ (208)
@@
$=\operatorname{Tr}\left[\hat{\rho}\hat{\Pi}(E)\right]^{2}.$ (209)
Inserting this inequality into Eq. (207), we get
$\left\lvert\int\overline{\mathrm{d}t}\ \operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert\leq\epsilon\sum_{E\in\mathcal{E}}\operatorname{Tr}\left[\hat{\rho}\hat{\Pi}(E)\right]=\epsilon,$ (210)
@@
$\sum_{E\in\mathcal{E}}\operatorname{Tr}\left[\hat{\rho}\hat{\Pi}(E)\right]=\operatorname{Tr}[\hat{\rho}]=1.$ (211)

### v2

@@ Appendix A Quantum Mazur-Suzuki inequality over finite times
Eq. (A.11) is nontrivial when the right hand side is larger than $A_{\mathrm{th}}^{2}$, which is the minimum value of the autocorrelator and corresponds to (global) thermalization. However, it only constrains thermalization over the same timescale (determined by $\Delta E$) in which the charge remains approximately conserved. For a noticeably broken conservation law (say, after a long timescale), $\delta Q_{k}^{2}[w_{2+}]$ is large, and consequently the second term within the absolute value that depends on $\delta Q_{k}^{2}[w_{2+}]$ weakens the bound, allowing the autocorrelator to decay further than with a perfect conservation law and potentially thermalize (as expected for systems with fewer conservation laws). For the question [54] of when this inequality can be saturated over finite timescales with a finite number of accessible charges $Q_{k}$, which are not necessarily energy projectors as in eigenstate thermalization [Eq. (12)], we note that this is possible when $\hat{A}]_{\Delta E}$ has most of its overlap with the $\hat{Q}_{k}$ rather than the $\hat{J}_{k}$ by Eq. (A.4) and the completeness relation Eq. (A.3) with $\hat{O}=[\hat{A}]_{\Delta E}$. It remains to be seen if this criterion can be used more systematically to determine when the inequality may be saturated in different classes of systems. It would also be interesting to consider the implications of the exact inequality in Eq. (A.11) for quantum transport problems, e.g., in a finite but large number of qubits.
@@ Appendix B Thermalization in dual-unitary quantum circuits
As a case study, mainly for illustrative purposes, let us consider the example of dual-unitary quantum circuits [55, 56, 57]. These are brickwork quantum circuits of, say, $N$ qubits (with Hilbert space dimension $D=2^{N}$), where the qubits are arranged in one dimension in a periodic chain and $2$-qubit unitary gates, each with the special property of “dual-unitarity”, simultaneously act on alternate pairs of qubits [^26] We consider qubits here for definiteness, but our conclusions should generalize to dual-unitary circuits for quddits [56, 57].. We will impose time translation invariance so that the resulting (Floquet) discrete-time system, with each step consisting of two alternating parallel applications of local gates, has well-defined (quasi-)energy levels. But we do not require spatial translation invariance, i.e., each $2$-qubit dual-unitary gate at a given time slice may be entirely different. The latter excludes spatial translation invariance as one of the special symmetries under which weak eigenstate thermalization may be directly shown for local observables in the infinite volume limit (assuming nondegenerate levels) [34, 35]; therefore, it is not yet clear if weak eigenstate thermalization is satisfied in general in our case. Even for translation-invariant circuits, it is not clear if the spectrum is necessarily nondegenerate for every single circuit of interest, which is required for thermalization to follow from eigenstate thermalization.
Under these circumstances, it can be shown that all autocorrelators of traceless single-site observables vanish for all time steps $\lvert t\rvert<N$, provided each unitary $2$-qubit gate satisfies the property of “dual unitarity”, for the details of which we refer to Ref. [55]. While the dynamics of specific (matrix product) initial states have been computed exactly for such circuits [92], and $f(E_{1},E_{2})$ in the ETH ansatz for certain observables has been estimated from autocorrelators [93] assuming the form in Eq. (3), our intention here is to show how more general conclusions on thermalization may be drawn in a much simpler way. Quantitatively, all traceless single-qubit observables $\widehat{\delta a}_{i}$ (with $i=1$ to $N$ indexing the qubit) satisfy:
$\operatorname{Tr}[\widehat{\delta a}_{i}(t)\widehat{\delta a}_{i}(0)]=0,\text{ for all }t:0<|t|<N.$ (B.1)
@@
$\left\lvert\langle\psi_{k}(t)\rvert\ \hat{a}_{i}\ \lvert\psi_{k}(t)\rangle-\frac{1}{2}\operatorname{Tr}_{i}[\hat{a}_{i}]\right\rvert<\epsilon\text{ for almost all }t\in[-T,T],\text{ for any }T>T_{0},$ (B.2)
@@
While we have expressed Eq. (B.2) for single-qubit observables, larger local observables on, say, $N_{S}$ consecutive qubits must also have vanishing autocorrelators except at extremely short $t=O(N_{S})$ and extremely long $t=\Omega(N)$ times, by our understanding of Ref. [55]. We therefore expect an analogue of Eq. (B.2) to hold for any traceless local observable spanning $N_{S}=o(N)$ consecutive sites.
Despite the strong thermalization behavior indicated by Eq. (B.2), it is known that some of these circuits show spectral correlations corresponding to dynamical non-ergodicity, e.g., Poisson spectral statistics [56]. These examples therefore also directly illustrate the logical separation between ergodic dynamics and the statistical mechanics of observables.
@@
$\hat{a}_{i}^{\text{co}}[t]=\hat{a}_{i+2t}.$ (B.3)
The dynamics of $\hat{a}_{i}^{\text{co}}$ in a dual unitary circuit is equivalent to the dynamics of the original $\hat{a}_{i}$ in a circuit formed by staggering each Floquet step of the dual unitary circuit with a (periodic) $2$-qubit shift/translation of the qubits in one dimension, in the opposite direction [94, 95]. In this case, however, the analogue of Eq. (B.1) is not satisfied for all dual unitary circuits, but only as an asymptotic equality (i.e. the autocorrelator is less than any given $\epsilon$) for special classes called mixing circuits at long but finite times (and for higher quddits, Bernoulli circuits at small times as well) [55, 56, 57]. Comoving single-particle observables can therefore thermalize in the sense of Eq. (B.2) for almost all initial states if and only if they are mixing or Bernoulli. While the “if” implication follows as above, the “only if” implication is due to Eq. (95) [combined with the cloning strategy of Sec. 7] which implies that if an observable thermalizes in almost all (computational) basis states, then its autocorrelator must thermalize.
@@ Appendix C Energy-band thermalization $\implies$ eigenspaces usually thermalize
$\Sigma_{d}=\Sigma_{d_{s}}\oplus\Sigma_{\delta d}$ (C.1)
@@
$n(\Sigma_{\delta d})\equiv\left(\sum_{\mathcal{H}_{\delta}(E)\in\Sigma_{\delta d}}1\right)<\frac{\epsilon^{2}}{\lambda^{2}}d.$ (C.2)
@@
Let us consider the implications of Eq. (C.2). First, we note that for almost all eigenspaces to thermalize, the ratio $n(\Sigma_{\delta d})/n(\Sigma_{d})$ should become negligible:
$\frac{n(\Sigma_{\delta d})}{n(\Sigma_{d})}\ll 1.$ (C.3)
In analogy with “weak eigenstate thermalization” [34, 35, 18], we will diagnose “weak eigenspace thermalization” if Eq. (C.3) is satisfied. By Proposition 3.3, this implies that thermalization o.a. occurs in arbitrary initial states supported on an energy shell $\Sigma_{d_{s}}$ that contains most of the eigenspaces in $\Sigma_{d}$. Determining weak eigenspace thermalization requires our accuracy for energy-band thermalization to be:
$\epsilon_{\text{weak}}\ll\lambda\ \sqrt{\frac{n(\Sigma_{d})}{d}}.$ (C.4)
@@
$\epsilon_{\text{weak}}>\frac{1}{N^{\alpha}},$ (C.5)
@@
$n(\Sigma_{d})>\frac{d}{N^{2\alpha}}.$ (C.6)
That we cannot show weak eigenspace thermalization with an accessible accuracy of energy-band thermalization without some constraint on the spectrum is why the title of this Appendix emphasizes the qualifier “usually”. Concretely, “usually” refers to systems satisfying Eq. (C.6); it implies that the spectrum must generally have a low degree of degeneracy. Nevertheless, we expect this condition to apply to a large class of systems, especially seeing that it allows $n(\Sigma_{d})\ll d$ (such as systems with a finite or at most weakly growing degree of degeneracy in each level, in the thermodynamic limit $d\to\infty$).
@@
$\epsilon<\frac{\lambda}{\sqrt{d}}$ (C.8)
in energy-band thermalization. This is consistent with Eq. (63). This constraint is independent of the energy spectrum, but as $\lambda<1$ for a nontrivial expression of eigenspace thermalization, Eq. (C.8) is not expected to be an accessible degree of accuracy (scaling as a power of $1/d$).
@@ D.1 Classical “eigenstate” thermalization without ergodicity
$=\sum_{k}\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}\pi_{\mathcal{P}_{k}}[\rho].$ (D.1)
@@
From Eq. (35), rewriting the left hand side in terms of the ergodic subsets $\mathcal{P}_{k}$ as in Eq. (D.1), we get
$\sum_{k}\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}\pi_{\mathcal{P}_{k}}[\rho_{A}]=\frac{\mu(A)}{\mu(\mathcal{P})}.$ (D.3)
@@
$\pi_{\mathcal{P}_{k}}[\rho_{A}]=\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(A)},$ (D.4)
using which Eq. (D.3) becomes (together with some manipulations on the right hand side using $\sum_{k}\mu(\mathcal{P}_{k})=\mu(\mathcal{P})$):
@@
$\implies\frac{1}{\mu(A)}\sum_{k}\mu(\mathcal{P}_{k})\frac{\mu(A\cap\mathcal{P}_{k})^{2}}{\mu(\mathcal{P}_{k})^{2}}$ $=\frac{1}{\mu(A)}\left[2\sum_{k}\mu(\mathcal{P}_{k})\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}\frac{\mu(A)}{\mu(\mathcal{P})}-\sum_{k}\mu(\mathcal{P}_{k})\frac{\mu(A)^{2}}{\mu(\mathcal{P})^{2}}\right].$ (D.5)
Collecting all the terms in Eq. (D.5), we get a weighted variance:
$\frac{1}{\mu(A)}\sum_{k}\mu(\mathcal{P}_{k})\left[\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}-\frac{\mu(A)}{\mu(\mathcal{P})}\right]^{2}=0.$ (D.6)
@@
$\frac{\mu(A\cap\mathcal{P}_{k})}{\mu(\mathcal{P}_{k})}-\frac{\mu(A)}{\mu(\mathcal{P})}=0.$ (D.7)
@@ D.2 Quantum eigenspace thermalization with degeneracies
D.2.1 Proposition 3.1: Quantum eigenstate thermalization implies thermalization o.a. given a nondegenerate spectrum [10, 15]
@@
$\left\lvert\int\overline{\mathrm{d}t}\ \operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert\leq\left(\sum_{n:\lvert E_{n}\rangle\in\Sigma_{d}}\langle E_{n}\rvert\hat{\rho}\lvert E_{n}\rangle\right)\max_{n:\lvert E_{n}\rangle\in\Sigma_{d}}\left\lvert\langle E_{n}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert$ (D.8)
@@
Rewriting Eq. (50) to include the $\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}$ term, we get:
@@
$=\sum_{E\in\mathcal{E}}\operatorname{Tr}\left[\hat{\Pi}(E)\hat{\rho}\hat{\Pi}(E)\hat{\Pi}(E)\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}(E)\right],$ (D.9)
@@
$\hat{\Pi}(E)=\hat{\Pi}(E)^{2}=\sum_{n:\lvert E_{n}\rangle\in\mathcal{H}_{E}}\lvert E_{n}\rangle\langle E_{n}\rvert$. (D.10)
Now applying the Cauchy-Schwarz inequality to each term in Eq. (D.9) with a given $E$, and the triangle inequality to consider each such term separately, we get
$\left\lvert\int\overline{\mathrm{d}t}\ \operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert$ $\leq\sum_{E\in\mathcal{E}}\sqrt{\operatorname{Tr}\left[\left\{\hat{\rho}\hat{\Pi}(E)\right\}^{2}\right]\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}(E)\right\}^{2}\right]}$ (D.11)
$\leq\epsilon\sum_{E\in\mathcal{E}}\sqrt{\operatorname{Tr}\left[\left\{\hat{\rho}\hat{\Pi}(E)\right\}^{2}\right]}.$ (D.12)
@@
$\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\rvert^{2}\leq\langle E_{n}\rvert\hat{\rho}\lvert E_{n}\rangle\langle E_{m}\rvert\hat{\rho}\lvert E_{m}\rangle,$ (D.13)
@@
$=\operatorname{Tr}\left[\hat{\rho}\hat{\Pi}(E)\right]^{2}.$ (D.14)
Inserting this inequality into Eq. (D.12), we get
$\left\lvert\int\overline{\mathrm{d}t}\ \operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert\leq\epsilon\sum_{E\in\mathcal{E}}\operatorname{Tr}\left[\hat{\rho}\hat{\Pi}(E)\right]=\epsilon,$ (D.15)
@@
$\sum_{E\in\mathcal{E}}\operatorname{Tr}\left[\hat{\rho}\hat{\Pi}(E)\right]=\operatorname{Tr}[\hat{\rho}]=1.$ (D.16)